# Studies of $Z$ boson decay into double $\Upsilon$ mesons at the NLO QCD accuracy

### v1

@@ I Introduction
In 2019, the CMS Collaboration measured the branching ratio of the $Z$ boson decaying to double $J/\psi$ CMS:2019wch , denoted as $\mathcal{B}_{Z\to J/\psi+J/\psi}<2.2\times 10^{-6}$. This upper limit was recently updated to be $1.4\times 10^{-6}$ by analyzing a larger sample of data CMS:2022fsq . At leading order (LO) in $\alpha_{s}$, the standard model provides a prediction of $\mathcal{B}_{Z\to J/\psi+J/\psi}\sim 10^{-12}$ Likhoded:2017jmx , which was upgraded to $10^{-10}$ by further evaluating the QED contributions resulting from the virtual-photon effects Gao:2022mwa . Recently, Li $et~{}al.$ suggested that the QCD corrections to the QCD and QED diagrams in $Z\to J/\psi+J/\psi$ would offset each another, maintaining the next-to-leading-order (NLO) results at the order of $10^{-10}$ z decay 40 .
@@
It is noteworthy that, in contrast to $J/\psi$, the substantial mass of the $b$ quark endows the $\Upsilon$ decay products with higher energy, rendering them more readily detectable. Additionally, the considerable mass of the $b$ quark typically leads to a more rapidly converging perturbative series in the expansion of $\alpha_{s}$ and $v^{2}$. Coupled with the significant production rates of Z bosons at the LHC, the proposed HL-LHC, or the future CEPC CEPC , the yields of double $\Upsilon$ mesons in $Z$ decay would offer an excellent laboratory for probing the mechanisms governing heavy quarkonium formation.
@@ II.1 Theoretical Framework
Figure 1: Representative QCD Feynman diagrams for $Z\to b\bar{b}[^{3}S_{1}^{1}]+b\bar{b}[^{3}S_{1}^{1},^{3}P_{J}^{1}]$. $\textbf{a}(\mathcal{M}_{\alpha^{\frac{1}{2}}\alpha_{s}})$ is the QCD tree-level diagram and $\textbf{b}-\textbf{f}(\mathcal{M}_{\alpha^{\frac{1}{2}}\alpha^{2}_{s}})$ are the NLO QCD corrections to $a$. Diagram b specifically represents the counter-term diagram.
[…]
The representative Feynman diagrams of $\mathcal{M}_{\alpha^{\frac{1}{2}}\alpha_{s}}$, $\mathcal{M}_{\alpha^{\frac{1}{2}}\alpha^{2}_{s}}$, $\mathcal{M}_{\alpha^{\frac{3}{2}}}$, and $\mathcal{M}_{\alpha^{\frac{3}{2}}\alpha_{s}}$ are displayed in Figs. 1 and 2. Fig. 1(a) ($\alpha^{\frac{1}{2}}\alpha_{s}$ order) demonstrates the QCD tree-level diagram (4 diagrams), whereas Figs. 1(b-f) elucidate the subsequent QCD NLO corrections, which comprise 56 individual one-loop diagrams and 20 counter-term diagrams. In Figs. 2(a,b) ($\alpha^{\frac{3}{2}}$ order) we observe the tree-level diagrams pertinent to QED, encompassing 8 diagrams in total. Furthermore, Figs. 2(c-k) depict the high-order corrections in $\alpha_{s}$, including 52 one-loop diagrams and 32 counter-term diagrams.
@@ III Phenomenological results
    Regarding the $b\bar{b}[^{3}S_{1}^{1}]$ yield associated with $b\bar{b}[^{3}P_{J}^{1}]$, the LO SDC $\hat{\Gamma}_{2,3}^{(0)}$ amplify $\hat{\Gamma}_{1}^{(0)}$ by approximately ten percents. With the inclusion of QCD corrections, $\hat{\Gamma}_{1}^{(1)}$ significantly enhances the $\hat{\Gamma}_{1}^{(0)}$ for the $b\bar{b}[^{3}P_{0,1}^{1}]$ process; however, the enhancement is milder for $b\bar{b}[^{3}P_{2}^{1}]$. In terms of QED processes, $\hat{\Gamma}_{3}^{(1)}$ slightly diminishes the LO SDC $\hat{\Gamma}_{3}^{(0)}$ for $b\bar{b}[^{3}P_{0,1}^{1}]$, while it substantially reduces that for $b\bar{b}[^{3}P_{2}^{1}]$. Consequently, due to the combined influence of enhancement and reduction effects, the QCD corrections ($\hat{\Gamma}_{2}^{(1)}$) have a substantial impact on $\hat{\Gamma}_{2}^{(0)}$ corresponding to $b\bar{b}[^{3}P_{0,1}^{1}]$, whereas the effect is more moderate for $b\bar{b}[^{3}P_{2}^{1}]$.
The ratio $\hat{\Gamma}^{\textrm{NLO}}_{\textrm{total}}/\hat{\Gamma}^{\textrm{LO}}_{\textrm{total}}$ suggests that QCD corrections can substantially enhance the LO results, underscoring the importance of our newly-calculated higher order terms in $\alpha_{s}$.
Figure 3: SDCs corresponding to $Z\to b\bar{b}[^{3}S_{1}^{1}]+b\bar{b}[^{3}S_{1}^{1},^{3}P_{J}^{1}]$ with respect to the renormalization scale. The $b$-quark mass is fixed at $m_{b}=4.7$ GeV. $\hat{\Gamma}_{\textrm{total}}$ denotes the collected sum of the contributions of $\hat{\Gamma}_{1,2,3}^{(0,1)}$.
@@
In Table 3, we confront our predictions with the CMS measurements. The subscripts “dir” and “fd” signify direct-production processes and feed-down effects, respectively. The ratio $\Gamma_{\textrm{fd}}/\Gamma_{\textrm{dir}}$ reveals that the feed-down contributions from the transitions of excited bottomonium states are crucial for $Z$ boson decay into a pair of $\Upsilon$ mesons. For example, in the case of $\Upsilon(1S)+\Upsilon(1S)$, feed-down contributions constitute approximately $30\%$ of the total results. Comparisons with experimental data indicate that the NLO predictions derived from the NRQCD framework are markedly below the upper threshold set by CMS, aligning with observations.

### v2

@@ I Introduction
In 2019, the CMS Collaboration measured the branching ratio of the $Z$ boson decaying to double $J/\psi$ CMS:2019wch , denoted as $\mathcal{B}_{Z\to J/\psi+J/\psi}<2.2\times 10^{-6}$. This upper limit was recently updated to be $1.4\times 10^{-6}$ by analyzing a larger sample of data CMS:2022fsq . At leading order (LO) in $\alpha_{s}$, the standard model provides a prediction of $\mathcal{B}_{Z\to J/\psi+J/\psi}\sim 10^{-12}$ Likhoded:2017jmx , which was upgraded to $10^{-10}$ by further evaluating the QED contributions resulting from the virtual-photon effects Gao:2022mwa . Recently, Li $et~{}al.$ suggested that the QCD corrections play a crucial role in enhancing the contributions from QCD diagrams and simultaneously diminishing the contributions from QED diagrams, ultimately maintaining the next-to-leading-order (NLO) results at the order of $10^{-10}$ z decay 40 .
@@
It is noteworthy that, the considerable mass of the $b$ quark typically leads to a more rapidly converging perturbative series in the expansion of $\alpha_{s}$ and $v^{2}$, compared to the case of the $c$ quark. Coupled with the significant production rates of Z bosons at the LHC, the proposed HL-LHC, or the future CEPC CEPC , the yields of double $\Upsilon$ mesons in $Z$ decay would offer an excellent laboratory for probing the mechanisms governing heavy quarkonium formation.
@@ II.1 Theoretical Framework
Figure 1: Representative QCD Feynman diagrams for $Z\to b\bar{b}[^{3}S_{1}^{1}]+b\bar{b}[^{3}S_{1}^{1},^{3}P_{J}^{1}]$. Diagram (a) (${\alpha^{\frac{1}{2}}\alpha_{s}}$ order) is the QCD tree-level diagram. Diagrams (b)-(f) $({\alpha^{\frac{1}{2}}\alpha^{2}_{s}}~{}\textrm{order})$ depict the NLO QCD corrections to (a). Diagram (b) specifically represents the counter-term diagram.
Figure 2: Representative QED Feynman diagrams for $Z\to b\bar{b}[^{3}S_{1}^{1}]+b\bar{b}[^{3}S_{1}^{1},^{3}P_{J}^{1}]$. Diagrams (a,b) $(\alpha^{\frac{3}{2}}$ order) are the QCD tree-level diagrams. Diagrams (c)-(k) ($\alpha^{\frac{3}{2}}\alpha_{s}$ order) depict the NLO QCD corrections to (a,b). Diagrams (c,g) specifically represent the counter-term diagram.
The representative Feynman diagrams of $\mathcal{M}_{\alpha^{\frac{1}{2}}\alpha_{s}}$, $\mathcal{M}_{\alpha^{\frac{1}{2}}\alpha^{2}_{s}}$, $\mathcal{M}_{\alpha^{\frac{3}{2}}}$, and $\mathcal{M}_{\alpha^{\frac{3}{2}}\alpha_{s}}$ are displayed in Figs. 1 and 2. Fig. 1(a) ($\alpha^{\frac{1}{2}}\alpha_{s}$ order) demonstrates the QCD tree-level diagram (4 diagrams), whereas Figs. 1(b)-(f) elucidate the subsequent QCD NLO corrections, which comprise 56 individual one-loop diagrams and 20 counter-term diagrams. In Figs. 2(a,b) ($\alpha^{\frac{3}{2}}$ order) we observe the tree-level diagrams pertinent to QED, encompassing 8 diagrams in total. Furthermore, Figs. 2(c)-(k) depict the high-order corrections in $\alpha_{s}$, including 52 one-loop diagrams and 32 counter-term diagrams.
@@ III Phenomenological results
    Regarding the $b\bar{b}[^{3}S_{1}^{1}]$ yield associated with $b\bar{b}[^{3}P_{J}^{1}]$, the LO SDCs $\hat{\Gamma}_{2,3}^{(0)}$ amplify $\hat{\Gamma}_{1}^{(0)}$ by approximately ten percents. With the inclusion of QCD corrections, $\hat{\Gamma}_{1}^{(1)}$ significantly enhances the $\hat{\Gamma}_{1}^{(0)}$ for the $b\bar{b}[^{3}P_{0,1}^{1}]$ process; however, the enhancement is milder for $b\bar{b}[^{3}P_{2}^{1}]$. In terms of QED processes, $\hat{\Gamma}_{3}^{(1)}$ slightly diminishes the LO SDC $\hat{\Gamma}_{3}^{(0)}$ for $b\bar{b}[^{3}P_{0,1}^{1}]$, while it substantially reduces that for $b\bar{b}[^{3}P_{2}^{1}]$. Consequently, due to the combined influence of enhancement and reduction effects, the QCD corrections ($\hat{\Gamma}_{2}^{(1)}$) have a substantial impact on $\hat{\Gamma}_{2}^{(0)}$ corresponding to $b\bar{b}[^{3}P_{0,1}^{1}]$, whereas the effect is more moderate for $b\bar{b}[^{3}P_{2}^{1}]$.
The ratio $\hat{\Gamma}^{\textrm{NLO}}_{\textrm{total}}/\hat{\Gamma}^{\textrm{LO}}_{\textrm{total}}$ in Tables 1 and 2 suggests that QCD corrections can substantially enhance the LO results, underscoring the importance of our newly-calculated higher order terms in $\alpha_{s}$.
Figure 3: SDCs corresponding to $Z\to b\bar{b}[^{3}S_{1}^{1}]+b\bar{b}[^{3}S_{1}^{1},^{3}P_{J}^{1}]$ with respect to the renormalization scale $\mu_{r}$. The $b$-quark mass is fixed at $m_{b}=4.7$ GeV. $\hat{\Gamma}_{\textrm{total}}$ denotes the collected sum of the contributions of $\hat{\Gamma}_{1,2,3}^{(0,1)}$.
@@
In Table 3, we confront our predictions with the CMS measurements. The subscripts “dir” and “fd” signify direct-production processes and feed-down effects, respectively. The ratio $\Gamma_{\textrm{fd}}/\Gamma_{\textrm{dir}}$ reveals that the feed-down contributions from the transitions of excited bottomonium states are crucial for $Z$ boson decay into a pair of $\Upsilon$ mesons. For example, in the case of $\Upsilon(1S)+\Upsilon(1S)$, feed-down contributions constitute approximately $30\%$ of the total results. Comparisons with experimental data indicate that the NLO predictions derived from the NRQCD framework are markedly below the upper threshold set by CMS.
@@
Note that, in $Z$ boson decay into double $\Upsilon$, besides the color-singlet processes of interest, color-octet (CO) processes are also possible. The primary CO contributions are derived from the process $Z\to b\bar{b}[^{3}S_{1}^{8}]+g^{*}$, followed by $g^{*}\to b\bar{b}[^{3}S_{1}^{8}]$. Given the ratio of $\frac{\langle{\mathcal{O}}^{H}(^{3}S_{1}^{8})\rangle}{\langle{\mathcal{O}}^{H}(^{3}S_{1}^{1})\rangle}\sim 10^{-3}$ Br3 , the branching ratio $\mathcal{B}_{Z\to\Upsilon(1S)+\Upsilon(1S)}$ due to CO contributions are estimated to be $10^{-14}$. Moreover, considering that experiments account for feed down from higher excited states, the semi-inclusive production of double $\Upsilon$ mesons in $Z$ boson decay, such as contributions from $b$(or $\bar{b}$) quark fragmentation into $\Upsilon$, should be considered. Utilizing the fragmentation function from Ref. z decay 4 , the branching ratio $\mathcal{B}_{Z\to\Upsilon(1S)+\Upsilon(1S)+X}$ is predicted to be $\sim 10^{-11}$, which is comparable with the results in Eq. (18).